Assume you are managing a food processing plant in Ethiopia. The demand function 

for one of your product is given as Qd=50-2p. (8 marks)

a) Find the point price elasticity if price is 15 ETB? Is it elastic or inelastic?

b) How do you interpret the elasticity result? 

c) In order to get more revenue what will be your recommendation. Is it to increase 

price or decrease price? Why? 

d) Describe at least four determinants of the price elasticity demand for the food 

product?


1
Expert's answer
2021-06-14T13:48:41-0400

Solution:

a.). Point price elasticity of demand (PEd) = "\frac{\%\triangle Qd}{\%\triangle P} = \frac{\triangle Qd}{\triangle P}\times \frac{P}{Q}"

Start by deriving Qd at the price of 15:

Qd = 50 – 2p

Qd = 50 – 2(15)

Qd = 50 – 30 = 20

Qd = 20

Calculate PEd:

Qd = 50 – 2p


"\frac{\triangle Qd}{\triangle P}" = The slope of the demand function = -2

Point price elasticity of demand (PEd) = "-2 \times \frac{15}{20} = -1.5"


The point price elasticity of demand (PEd) of 1.5 is elastic.

 

b.). The point price elasticity of demand is greater than 1 which means that the demand of the product is elastic. This can be interpreted as consumers being very sensitive to changes in price. Therefore, a slight percentage increase in price will lead to a larger percentage decrease in quantity demanded and vice versa.

 

c.). The best alternative will be to reduce the price of the product, in order to increase the product's revenue. When the demand is elastic at a certain price level, the company should reduce its price since the percentage drop in price will result in an even larger percentage increase in the quantity sold, hence increasing total revenue.

 

d.). The four determinants of the price elasticity of demand for the food product are listed below:

1.). Whether the good is a luxury or necessity – Luxury products tend to be highly elastic since a fall in their prices will lead to an increase in demand compared to necessity items such as essential goods which are inelastic, since demand doesn’t change much with price changes.

 

2.). Availability of close substitutes – The price elasticity of demand would be considered elastic when consumers can easily substitute the food product with another readily available product, which they regard as similar. On the other hand, if consumers cannot substitute the food product, then the product will experience inelastic demand.

 

3.). Income levels – Rich people who earn high incomes have a price elasticity of demand that is lower compared to poor people who are low-income earners. This is because the wealthy, who are high-income earners are not affected much by changes in prices compared to the poor, who are low-income earners who are highly affected by changes in the prices of a product.

 

4.). Price levels – Products that have high prices tend to be demand elastic, while products which have lower prices are demand inelastic since their demand doesn’t change much with changes in price.